OpenStudy (anonymous):

Let G = { a + bi | I a + bi I ≤ 4 } Which of the following are in G. (a) 3 + 3i (b) 3i (c) -4 + 3i (d) i^5 (e) 4 - 4i (f) 6 + 2i

OpenStudy (anonymous):

I swear I would not be asking if I knew how to do it. I'm not being lazy I have no clue :( My teacher is lazy

OpenStudy (kinggeorge):

Well, to find out, all you have to do is find if\[\sqrt{a^2+b^2} \leq 4\]In other words, \[a^2+b^2 \leq 16\]

OpenStudy (kinggeorge):

Also, \(i^5=i\) if you didn't know.

OpenStudy (kinggeorge):

By my calculations, b and d should be in G, but the rest aren't.

OpenStudy (anonymous):

Thank you very much but how do you find out what should be in there

OpenStudy (kinggeorge):

By the formula I posted above. All the numbers are of the form \(a+bi\) for some \(a, b \in \mathbb{R} \) so then you just use the formula I wrote above.

OpenStudy (anonymous):

Hmm shouldn't that mean a should be in there as well 3 + 3i = 3 + √-9 ≤ 16?

OpenStudy (kinggeorge):

No, since \(a=b=3\) so \(3^2+3^2=18\) and this is not less than 16.

OpenStudy (kinggeorge):

Remember you need to square the numbers and then add them.

OpenStudy (anonymous):

Right, I'm an idiot lol thank you very much

OpenStudy (kinggeorge):

You're welcome.